# Pricing Options on Forwards in Function-Valued Affine Stochastic Volatility Models

### v1

@@ 4.4. Pricing in the operator-valued Barndorff–Nielsen–Shephard stochastic volatility model
$\mathop{}\!\textnormal{e}^{(t-s)\mathcal{L}_{Q}}(f_{n}\otimes f_{j})=\mathop{}\!\textnormal{e}^{-(a_{n}+a_{j})(t-s)}\,f_{n}\otimes f_{j}.$
@@
$D^{1/2}\psi_{1}(s)=\frac{\nu+i\lambda}{\sqrt{2}}\sum_{n\geq 1}\frac{1}{n}\,f_{n}(s+\vartheta)\,f_{n},\qquad D^{1/2}\Upsilon^{\vartheta}\;=\;\frac{1}{\sqrt{2}}\sum_{j\geq 1}\frac{1}{j}\,c_{j}(\vartheta)\,f_{j},\quad c_{j}(\vartheta):=\langle f_{j},S^{*}(\vartheta)h_{0}\rangle_{w}.$
@@
$\psi_{2}(t)f_{n}$ $=-\!\sum_{j\geq 1}\!\int_{0}^{t}e^{-(a_{n}+a_{j})(t-s)}\Big{[}\frac{(\nu+i\lambda)^{2}}{4\,nj}\,f_{n}(s+\vartheta)f_{j}(s+\vartheta)+\frac{(\nu+i\lambda)}{2\,nj}\,c_{j}(\vartheta)\,f_{n}(s+\vartheta)\Big{]}\mathop{}\!\textnormal{d}s\;f_{j},$
@@
$\langle f_{n},\psi_{2}(t)f_{n}\rangle_{w}=-\int_{0}^{t}e^{-2a_{n}(t-s)}\left[\frac{(\nu+i\lambda)^{2}}{4\,n^{2}}\,f_{n}(s+\vartheta)^{2}+\frac{(\nu+i\lambda)}{2\,n^{2}}\,c_{n}(\vartheta)\,f_{n}(s+\vartheta)\right]\mathop{}\!\textnormal{d}s,$
from which we deduce formula (4.15). Finally, since for the compound-Poisson input without finite-variation drift it holds that $\partial_{t}\phi(t,u)=F(\psi_{2}(t,u))=\beta\big{(}e^{-\langle D,\psi_{2}(t,u)\rangle}-1\big{)}$, integrating over $[0,t]$ yields (4.14). ∎
@@
$\mathbb{E}_{\mathbb{Q}}\big{[}(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big{]}$ $=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\!\Big{(}-\phi(t,u)+\langle X_{0},\psi_{1}(t,u)\rangle_{w}-\langle Y_{0},\psi_{2}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda$ (4.17)
$=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\!\Big{(}-\beta\!\int_{0}^{t}\!\big{(}\mathop{}\!\textnormal{e}^{-\langle D,\psi_{2}(s,u)\rangle}-1\big{)}\,\mathop{}\!\textnormal{d}s\big{)}$
$\qquad\qquad\times\exp\!\Big{(}(\nu+i\lambda)\,X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda,$
where $u_{1}=(\nu+i\lambda)u_{\vartheta}$ with $\nu>1$, and $\langle D,\psi_{2}(t,u)\rangle$ is as in (4.15), with $Q=-\sum_{n\geq 1}a_{n}\,f_{n}\otimes f_{n}$ ($a_{n}\geq 0$) and $c_{n}(\vartheta):=\langle f_{n},S^{*}(\vartheta)h_{0}\rangle_{w}$. The term $\langle X_{0},\psi_{1}(t,u)\rangle_{w}=(\nu+i\lambda)X_{0}(t+\vartheta)$ comes from $\psi_{1}(t,u)=S^{*}(t)u_{1}$.
@@
$\mathbb{E}_{\mathbb{Q}}\big{[}(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big{]}$ $\approx\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big{(}-\beta\!\int_{0}^{t}\!\big{(}e^{-\langle D,\psi_{2}^{(N)}(s,u)\rangle}-1\big{)}\,\mathop{}\!\textnormal{d}s\Big{)}$ (4.18)
$\qquad\qquad\times\exp\Big{(}(\nu+i\lambda)\,X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}^{(N)}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda,$
@@
$\langle D,\psi_{2}^{(N)}(t,u)\rangle=-\int_{0}^{t}\Big{[}\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{N}\frac{1}{n^{4}}f_{n}(s+\vartheta)^{2}+\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{N}\frac{c_{n}(\vartheta)}{n^{4}}f_{n}(s+\vartheta)\Big{]}\mathop{}\!\textnormal{e}^{-2a_{n}(t-s)}\,\mathop{}\!\textnormal{d}s.$
In our experiments, we specialize to $a_{n}=\frac{1}{2n^{2}}$, then $e^{-2a_{n}(t-s)}=\mathop{}\!\textnormal{e}^{-(t-s)/n^{2}}$.
@@
Figure 3. Monte Carlo v.s. analytical formula, with $Y_{0}=D,h_{0}=1,\alpha=0.1,T_{0}=1,N=5$, strike $K=1$ and the intensity $\beta=1$.
@@ 4.5. Running time analysis
Table 5 presents the runtime comparison for option pricing between the affine semi-analytical method and the Monte Carlo simulation, based on the experiments (Monte Carlo v.s. analytical formula) in Section 4.3 and 4.4. Both experiments were conducted on a machine equipped with an AMD Ryzen 5 PRO 7545U CPU (6 cores, 3.20 GHz base clock), 32 GB DDR5 RAM. Algorithms were implemented in Python 3.12.8 using NumPy 1.26.4 and SciPy 1.14.1. Each reported runtime is the average over 10 independent runs, measured with the "time" module. The results clearly show that the affine semi-analytical method is approximately 6000 times faster than the Monte Carlo simulation, which indicates its significant advantage for practical applications where computational efficiency is critical.
@@
Lévy driven model 0.03 173.70
BNS model 0.03 181.79
Table 5. Runtime (seconds) comparison: Affine formula v.s. Monte Carlo (1 million scenarios). On the option pricing for pure-jump Lévy driven model (Section 4.3) and the BNS model (Section 4.4)
@@ 5.1. Real and complex moments for the affine pure-jump stochastic volatility model
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and let $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ be the generator of a strongly continuous semigroup, let $D\in\mathcal{L}(H)$ be a positive self-adjoint operator, and let $\Upsilon\in H$. Then for every $u=(u_{1},u_{2})\in H\times\mathcal{O}$ there exist a positive real number $T_{\sf q_{2}}$ and a unique solution $({\sf p}(\cdot,u),{\sf q}_{1}(\cdot,u),+{\sf q}_{2}(\cdot,u))$ to (2.7a)-(2.7c) on $0,T_{\sf q_{2}})$.
@@
$\qquad-\int_{\mathcal{H}^{+}\setminus\{0\}}\big{(}\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}-1\big{)}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}.$
@@
Observe that the function $\mathcal{R}_{u_{1}}(t,\cdot)$ is locally Lipschitz continuous on $\mathcal{O}$ for every $t\geq 0$ and $u_{1}\in H$. Since $\mathcal{O}$ is assumed to be open, it follows from standard ODE results (see e.g. [36, Chapter 6, Proposition 1.2]) that for every $u_{2}\in\mathcal{O}$ there exist a $T_{\sf q_{2}}>0$ and a unique solution ${\sf q}_{2}(\cdot,u)$ of (5.1) on $0,T_{{\sf q}_{2}})$ with
@@
To compute exponential moments for the process $(X_{t},\mathsf{Y}_{t})_{t\geq 0}$ in (2.4), we first consider the joint process $(X^{(n)}_{t},\mathsf{Y}_{t})_{t\geq 0}$ obtained by replacing $\mathcal{A}$ in (2.4) by its Yosida approximation $\mathcal{A}^{(n)}:=n\mathcal{A}(nI-\mathcal{A})^{-1}$. Namely, we consider the process $X^{(n)}\colon[0,\infty)\times\Omega\rightarrow H$ given by the solution to (see [21, Proposition 6.4] for the existence of the solution to $X^{(n)}$)
$X^{(n)}_{t}=x+\int_{0}^{t}(\mathcal{A}^{(n)}X^{(n)}_{s}+D^{1/2}X^{(n)}_{s}D^{1/2}\Upsilon)\,\mathop{}\!\textnormal{d}s+\int_{0}^{t}D^{1/2}X_{s}^{1/2}\,\mathop{}\!\textnormal{d}W_{s}\,,\quad t\geq 0.$ (5.2)
The use of the approximation allows us to exploit the semimartingale theory and to apply the Itô formula and standard techniques in order to compute exponential moments for $(X^{(n)}_{t},\mathsf{Y}_{t})_{t\geq 0}$. Then we take the limit when $n$ goes to $\infty$. Observe that from [21, Proposition 7.5] it holds
@@
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, let $(X,Y)$ be the stochastic volatility model (2.4), let $u=(u_{1},u_{2})\in H\times\mathcal{O}$, and let $({\sf P}(\cdot,u),{\sf q}_{1}(\cdot),u),{\sf q}_{2}(\cdot,u))$ be the mild solution to the Riccati equation (2.7) up to time $T$ the existence of which is guaranteed by Proposition 5.2. Then we have:
$\mathbb{E}\left[\mathop{}\!\textnormal{e}^{\langle X_{T},u_{1}\rangle_{H}-\langle{\sf Y}_{T},u_{2}\rangle}\right]<\infty.$ (5.4)
@@
$\mathbb{E}\left[\mathop{}\!\textnormal{e}^{\langle X_{T},u_{1}\rangle_{H}-\langle{\sf Y}_{T},u_{2}\rangle}\mid\mathcal{F}_{t}\right]=\mathop{}\!\textnormal{e}^{-{\sf P}(T-t,u)+\langle X_{t},{\sf q}_{1}(T-t,u)\rangle_{H}-\langle{\sf Y}_{t},{\sf q}_{2}(T-t,u)\rangle}.$ (5.5)
@@
$g_{u}^{(n)}(t,x,y)=\exp{\{-{\sf P}^{(n)}(T-t,u)+\langle x,{\sf q}^{(n)}_{1}(T-t,u)\rangle_{H}-\langle y,{\sf q}^{(n)}_{2}(T-t,u)\rangle\}}\,.$
Similarly to the proof of [16, Theorem 3.3], we apply the Itô’s formula to $g_{u}^{(n)}(t,X^{(n)}_{t},{\sf Y}_{t})$ to deduce that it is a local martingale.
@@
$\mathbb{E}\left[\mathop{}\!\textnormal{e}^{\langle X^{(n)}_{T},u\rangle_{H}-\langle{\sf Y}_{T},u\rangle}\right]$ $=\mathbb{E}\left[g_{u}^{(n)}(T,X_{T}^{(n)},{\sf Y}_{T})\right]$
@@
$\mathbb{E}\left[\mathop{}\!\textnormal{e}^{\langle X_{T}^{(n)},u_{1}\rangle_{H}-\langle{\sf Y}_{T},u_{2}\rangle}\mid\mathcal{F}_{t}\right]$ $=\mathop{}\!\textnormal{e}^{-{\sf P}^{(n)}(T-t,u)+\langle X_{t},{\sf q}_{1}^{(n)}(T-t,u)\rangle_{H}-\langle{\sf Y}_{t},{\sf q}_{2}^{(n)}(T-t,u)\rangle}.$
Following similar arguments as in the proof of [16, Proposition 3.5], we deduce that
@@
$\lim_{n\rightarrow\infty}\left(\sup_{t\in[0,T]}\|{\sf q}_{1}^{(n)}(t,u)-{\sf q}_{1}(t,u)\|_{H}+\sup_{t\in[0,T]}\|{\sf q}_{2}^{(n)}(t,u)-{\sf q}_{2}(t,u)\|_{2}\right)\,.$

### v2

@@ 4.4. Pricing in the operator-valued Barndorff–Nielsen–Shephard stochastic volatility model
$\mathop{}\!\mathrm{e}^{(t-s)\mathcal{L}_{Q}}(f_{n}\otimes f_{j})=\mathop{}\!\mathrm{e}^{-(a_{n}+a_{j})(t-s)}\,f_{n}\otimes f_{j}.$
@@
$D^{1/2}\psi_{1}(s)=\frac{\nu+i\lambda}{\sqrt{2}}\sum_{n\geq 1}\frac{1}{n}\,f_{n}(s+\vartheta)\,f_{n},\qquad D^{1/2}\Upsilon^{\vartheta}\;=\;\frac{1}{\sqrt{2}}\sum_{j\geq 1}\frac{1}{j}\,c_{j}(\vartheta)\,f_{j},\quad c_{j}(\vartheta):=-\tfrac{1}{2}\langle f_{j},S^{*}(\vartheta)h_{0}\rangle_{w}.$
@@
$\psi_{2}(t)f_{n}$ $=-\!\sum_{j\geq 1}\!\int_{0}^{t}\mathop{}\!\mathrm{e}^{-(a_{n}+a_{j})(t-s)}\Big[\frac{(\nu+i\lambda)^{2}}{4\,nj}\,f_{n}(s+\vartheta)f_{j}(s+\vartheta)+\frac{(\nu+i\lambda)}{2\,nj}\,c_{j}(\vartheta)\,f_{n}(s+\vartheta)\Big]\mathop{}\!\mathrm{d}s\;f_{j},$
@@
$\langle f_{n},\psi_{2}(t)f_{n}\rangle_{w}=-\int_{0}^{t}\mathop{}\!\mathrm{e}^{-2a_{n}(t-s)}\left[\frac{(\nu+i\lambda)^{2}}{4\,n^{2}}\,f_{n}(s+\vartheta)^{2}+\frac{(\nu+i\lambda)}{2\,n^{2}}\,c_{n}(\vartheta)\,f_{n}(s+\vartheta)\right]\mathop{}\!\mathrm{d}s,$
from which we deduce formula (4.15). Finally, since for the compound-Poisson input without finite-variation drift it holds that $\partial_{t}\phi(t,u)=F(\psi_{2}(t,u))=-\beta\big(\mathop{}\!\mathrm{e}^{-\langle D,\psi_{2}(t,u)\rangle}-1\big)$, integrating over $[0,t]$ yields (4.14). ∎
@@
$\mathbb{E}_{\mathbb{Q}}\big[(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big]$ $=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big(-\phi(t,u)+\langle X_{0},\psi_{1}(t,u)\rangle_{w}-\langle Y_{0},\psi_{2}(t,u)\rangle\Big)\,\mathop{}\!\mathrm{d}\lambda$ (4.17)
$=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big(\beta\!\int_{0}^{t}\!\big(\mathop{}\!\mathrm{e}^{-\langle D,\psi_{2}(s,u)\rangle}-1\big)\,\mathop{}\!\mathrm{d}s\big)$
$\qquad\qquad\times\exp\Big((\nu+i\lambda)\,X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}(t,u)\rangle\Big)\,\mathop{}\!\mathrm{d}\lambda,$
where $u_{1}=(\nu+i\lambda)u_{\vartheta}$ with $\nu>1$, and $\langle D,\psi_{2}(t,u)\rangle$ is as in (4.15), with $Q=-\sum_{n\geq 1}a_{n}\,f_{n}\otimes f_{n}$ ($a_{n}\geq 0$) and $c_{n}(\vartheta):=-\tfrac{1}{2}\langle f_{n},S^{*}(\vartheta)h_{0}\rangle_{w}$. The term $\langle X_{0},\psi_{1}(t,u)\rangle_{w}=(\nu+i\lambda)X_{0}(t+\vartheta)$ comes from $\psi_{1}(t,u)=S^{*}(t)u_{1}$.
@@
$\mathbb{E}_{\mathbb{Q}}\big[(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big]$ $\approx\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big(\beta\!\int_{0}^{t}\!\big(\mathop{}\!\mathrm{e}^{-\langle D,\psi_{2}^{(N)}(s,u)\rangle}-1\big)\,\mathop{}\!\mathrm{d}s\Big)$ (4.18)
$\qquad\qquad\times\exp\Big((\nu+i\lambda)\,X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}^{(N)}(t,u)\rangle\Big)\,\mathop{}\!\mathrm{d}\lambda,$
@@
$\langle D,\psi_{2}^{(N)}(t,u)\rangle=-\int_{0}^{t}\Big[\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{N}\frac{1}{n^{4}}f_{n}(s+\vartheta)^{2}+\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{N}\frac{c_{n}(\vartheta)}{n^{4}}f_{n}(s+\vartheta)\Big]\mathop{}\!\mathrm{e}^{-2a_{n}(t-s)}\,\mathop{}\!\mathrm{d}s.$
In our experiments, we specialize to $a_{n}=\frac{1}{2n^{2}}$, then $\mathop{}\!\mathrm{e}^{-2a_{n}(t-s)}=\mathop{}\!\mathrm{e}^{-(t-s)/n^{2}}$.
@@
Figure 4. Conditional Monte Carlo vs. affine formula, with $Y_{0}=D,h_{0}=1,\alpha=0.1,T_{0}=1,N=5$, strike $K=1$ and the intensity $\beta=1$.
4.5. Pricing in the affine jump model with state-dependent intensity
As a fourth example we validate the affine semi-explicit formula for the model from Section 2.1 in the case of a state-dependent compound-Poisson jump kernel. The covariance process is now an operator-valued OU process driven by jumps whose arrival rate depends on the current state.
Model specification.
We take $N=5$ basis functions $(f_{n})_{n=1}^{5}$ as in (4.2) and set $D=\frac{1}{2}\sum_{n=1}^{N}n^{-2}f_{n}\otimes f_{n}$. The covariance process $\mathsf{Y}_{t}=\sum_{n=1}^{N}y_{n}(t)\,f_{n}\otimes f_{n}$ satisfies
$\mathop{}\!\mathrm{d}\mathsf{Y}_{t}=-(Q\mathsf{Y}_{t}+\mathsf{Y}_{t}Q)\,\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}J_{t},\qquad\mathsf{Y}_{0}=\sum_{n=1}^{N}\tfrac{1}{2n^{2}}\,f_{n}\otimes f_{n},$ (4.19)
with $Q=\sum_{n=1}^{N}\kappa_{n}\,f_{n}\otimes f_{n}$, $\kappa_{n}=n^{-2}$, so each diagonal entry satisfies $\mathop{}\!\mathrm{d}y_{n}=-2\kappa_{n}y_{n}\,\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}J_{n}$. The jump process $J_{t}$ has a single jump direction $\Xi=\sum_{n=1}^{N}\xi_{n}\,f_{n}\otimes f_{n}$, $\xi_{n}=(2n^{2})^{-1}$, and state-dependent intensity
$\lambda(\mathsf{Y})=\lambda_{0}+\sum_{n=1}^{N}\frac{c_{n}}{\|\xi\|_{2}^{2}}\,y_{n},\qquad\lambda_{0}=0.05,\quad c_{n}=\frac{0.05}{n^{2}},\quad\|\xi\|_{2}^{2}=\sum_{n=1}^{N}\xi_{n}^{2}.$ (4.20)
The forward dynamics (4.3) and HJM drift $\Upsilon^{\vartheta}=-\frac{1}{2}S^{*}(\vartheta)h_{0}$ are as in the previous examples.
Numerical Riccati scheme.
Unlike the pure-jump Lévy case (Lemma 4.1), no closed-form Riccati solution is available due to the state-dependent jump intensity. In the single-direction diagonal specification implemented below, the generalized Riccati equations from Theorem 2.5 reduce, for the matrix coefficients $q_{ij}(\tau)=\langle f_{i},\psi_{2}(\tau,u)f_{j}\rangle_{w}$, to the system
$\partial_{\tau}q_{ij}=-(\kappa_{i}+\kappa_{j})\,q_{ij}-\frac{c^{2}}{4\,ij}\,f_{i}(\vartheta+\tau)\,f_{j}(\vartheta+\tau)+\frac{c}{4\,ij}\,f_{i}(\vartheta+\tau)\,f_{j}(\vartheta)-\delta_{ij}\,\frac{c_{i}}{\|\xi\|_{2}^{2}}\!\left(\exp\Bigl(-\sum_{n=1}^{N}\xi_{n}\,q_{nn}\Bigr)-1\right),$ (4.21)
where $c=\nu+i\lambda$. We solve (4.21) numerically by a forward Euler scheme with step $\delta=1/365$, and recover option prices via the Fourier inversion formula (3.2) with truncation $N_{\rm trunc}=10$.
Monte Carlo benchmark.
Given a simulated covariance path $(y_{n}(t_{l}))$, obtained by Euler–Maruyama with state-dependent Poisson arrivals and step $\delta=1/365$, the log-forward $\log F(T,T+\vartheta)$ is conditionally Gaussian with conditional mean and variance
$\mu_{\rm cond}$ $=X_{0}(T+\vartheta)-\frac{\delta}{4}\sum_{l}\sum_{n=1}^{N}\frac{y_{n}(t_{l})}{n^{2}}\,f_{n}(\vartheta)\,f_{n}(\vartheta+s_{l}),$
$\sigma_{\rm cond}^{2}$ $=\frac{\delta}{2}\sum_{l}\sum_{n=1}^{N}\frac{y_{n}(t_{l})}{n^{2}}\,f_{n}(\vartheta+s_{l})^{2},$
where $s_{l}=(l+\frac{1}{2})\delta$. Option prices are obtained by averaging the Black formula over simulated paths. We use $\alpha=0.1$, $h_{0}=1$, $T=1$, $K=1$, and $N_{\rm sim}\in\{1\mathrm{k},10\mathrm{k},50\mathrm{k},100\mathrm{k},200\mathrm{k}\}$.
Figure 5 shows the option prices as a function of time to maturity $\vartheta$. The Monte Carlo estimates converge to the semi-explicit affine prices as $N_{\rm sim}$ increases, confirming consistency of the numerical Riccati scheme with direct simulation of the state-dependent model.
Figure 5. Euler conditional Monte Carlo vs. affine semi-explicit formula for the state-dependent jump model, with $\alpha=0.1$, $h_{0}=1$, $T=1$, $K=1$, $N=5$, $\lambda_{0}=0.05$, $\kappa_{n}=n^{-2}$, $\xi_{n}=(2n^{2})^{-1}$.
@@ 4.6. Running time analysis
We report runtimes from the canonical root benchmark entry point python main_new.py runtime_benchmarks_paper, which stores the corresponding machine metadata and timings in runtime_benchmarks_paper.json. The benchmark was run on a Lenovo ThinkPad X1 Carbon Gen 11 (model 21HM006WMH) with a 13th Gen Intel Core i7-1355U processor and 32 GB RAM, under Python 3.12.2, NumPy 2.2.5, and SciPy 1.15.2. For the explicit Lévy and BNS models we use 100 000100\,000 conditional Monte Carlo paths, while for the state-dependent model we use 200 000200\,000 paths. Each entry in Table 5 is a single wall-clock evaluation of the corresponding pricing routine in the canonical root implementation.
@@
Lévy driven model 0.16 88.89
BNS model 0.08 103.51
State-dependent jump model 55.01 71.78
Table 5. Canonical runtime benchmark generated by python main_new.py runtime_benchmarks_paper on the Lenovo ThinkPad X1 Carbon Gen 11 described above. The Lévy and BNS rows use 100 000100\,000 Monte Carlo scenarios; the state-dependent row uses 200 000200\,000 scenarios.
These timings confirm the qualitative picture already seen in the pricing figures: for the explicit Lévy and BNS specifications, the affine formulas are orders of magnitude faster than the corresponding Monte Carlo benchmark, whereas in the state-dependent example the runtime gap is much smaller, due to the complex generalized Riccati equations which do not have closed form solutions beyond the OU case. Developing fast schemes for this type of infinite-dimensional Riccati equations is a promising research question left for the future.
@@ 5.1. Real and complex moments for the affine pure-jump stochastic volatility model
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and let $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ be the generator of a strongly continuous semigroup, let $D\in\mathcal{L}(H)$ be a positive self-adjoint operator, and let $\Upsilon\in H$. Then for every $u=(u_{1},u_{2})\in H\times\mathcal{O}$ there exist a positive real number $T_{\sf q_{2}}$ and a unique solution $({\sf p}(\cdot,u),{\sf q}_{1}(\cdot,u),{\sf q}_{2}(\cdot,u))$ to (2.7a)-(2.7c) on $0,T_{\sf q_{2}})$.
@@
$\qquad-\int_{\mathcal{H}^{+}\setminus\{0\}}\big(\mathop{}\!\mathrm{e}^{-\langle\xi,u\rangle}-1\big)\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}}.$
@@
Observe that the function $\mathcal{R}_{u_{1}}(t,\cdot)$ is locally Lipschitz continuous on $\mathcal{O}$ for every $t\geq 0$ and $u_{1}\in H$. Since $\mathcal{O}$ is assumed to be open, it follows from standard ODE results (see e.g. [34, Chapter 6, Proposition 1.2]) that for every $u_{2}\in\mathcal{O}$ there exist a $T_{\sf q_{2}}>0$ and a unique solution ${\sf q}_{2}(\cdot,u)$ of (5.1) on $0,T_{{\sf q}_{2}})$ with
@@
To compute exponential moments for the process $(X_{t},\mathsf{Y}_{t})_{t\geq 0}$ in (2.4), we first consider the joint process $(X^{(n)}_{t},\mathsf{Y}_{t})_{t\geq 0}$ obtained by replacing $\mathcal{A}$ in (2.4) by its Yosida approximation $\mathcal{A}^{(n)}:=n\mathcal{A}(nI-\mathcal{A})^{-1}$. Namely, we consider the process $X^{(n)}\colon[0,\infty)\times\Omega\rightarrow H$ given by the solution to (see [19, Proposition 6.4] for the existence of the solution to $X^{(n)}$)
$X^{(n)}_{t}=x+\int_{0}^{t}(\mathcal{A}^{(n)}X^{(n)}_{s}+D^{1/2}\mathsf{Y}_{s}D^{1/2}\Upsilon)\,\mathop{}\!\mathrm{d}s+\int_{0}^{t}D^{1/2}\mathsf{Y}_{s}^{1/2}\,\mathop{}\!\mathrm{d}W_{s}\,,\quad t\geq 0.$ (5.2)
The use of the approximation allows us to exploit the semimartingale theory and to apply the Itô formula and standard techniques in order to compute exponential moments for $(X^{(n)}_{t},\mathsf{Y}_{t})_{t\geq 0}$. Then we take the limit when $n$ goes to $\infty$. Observe that from [19, Proposition 7.5] it holds
@@
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, let $(X,Y)$ be the stochastic volatility model (2.4), let $u=(u_{1},u_{2})\in H\times\mathcal{O}$, and let $({\sf P}(\cdot,u),{\sf q}_{1}(\cdot,u),{\sf q}_{2}(\cdot,u))$ be the mild solution to the Riccati equation (2.7) up to time $T$ the existence of which is guaranteed by Proposition 5.2. Then we have:
$\mathbb{E}\left[\mathop{}\!\mathrm{e}^{\langle X_{T},u_{1}\rangle_{H}-\langle{\sf Y}_{T},u_{2}\rangle}\right]<\infty.$ (5.4)
@@
$\mathbb{E}\left[\mathop{}\!\mathrm{e}^{\langle X_{T},u_{1}\rangle_{H}-\langle{\sf Y}_{T},u_{2}\rangle}\mid\mathcal{F}_{t}\right]=\mathop{}\!\mathrm{e}^{-{\sf P}(T-t,u)+\langle X_{t},{\sf q}_{1}(T-t,u)\rangle_{H}-\langle{\sf Y}_{t},{\sf q}_{2}(T-t,u)\rangle}.$ (5.5)
@@
$g_{u}^{(n)}(t,x,y)=\exp\left\{-{\sf P}^{(n)}(T-t,u)+\langle x,{\sf q}^{(n)}_{1}(T-t,u)\rangle_{H}-\langle y,{\sf q}^{(n)}_{2}(T-t,u)\rangle\right\}\,.$
Similarly to the proof of [15, Theorem 3.3], we apply the Itô’s formula to $g_{u}^{(n)}(t,X^{(n)}_{t},{\sf Y}_{t})$ to deduce that it is a local martingale.
@@
$\mathbb{E}\left[\mathop{}\!\mathrm{e}^{\langle X^{(n)}_{T},u\rangle_{H}-\langle{\sf Y}_{T},u\rangle}\right]$ $=\mathbb{E}\left[g_{u}^{(n)}(T,X_{T}^{(n)},{\sf Y}_{T})\right]$
@@
$\mathbb{E}\left[\mathop{}\!\mathrm{e}^{\langle X_{T}^{(n)},u_{1}\rangle_{H}-\langle{\sf Y}_{T},u_{2}\rangle}\mid\mathcal{F}_{t}\right]$ $=\mathop{}\!\mathrm{e}^{-{\sf P}^{(n)}(T-t,u)+\langle X_{t},{\sf q}_{1}^{(n)}(T-t,u)\rangle_{H}-\langle{\sf Y}_{t},{\sf q}_{2}^{(n)}(T-t,u)\rangle}.$
Following similar arguments as in the proof of [15, Proposition 3.5], we deduce that
@@
$\lim_{n\rightarrow\infty}\left(\sup_{t\in[0,T]}\|{\sf q}_{1}^{(n)}(t,u)-{\sf q}_{1}(t,u)\|_{H}+\sup_{t\in[0,T]}\|{\sf q}_{2}^{(n)}(t,u)-{\sf q}_{2}(t,u)\|_{2}\right)=0\,.$